How many moles of potassium chloride are there in 1 liter of a 5.5 molar solution?
1
Expert's answer
2020-05-13T14:20:37-0400

Solution:

"c(KCl)=\frac{n(KCl)}{V(KCl)} \\n(KCl)=c(KCl)*V(KCl)=5.5*1=5.5(mol)"

Answer:

n(KCl)=5.5 mol